OpenStudy (anonymous):

Consider f(n) = lg(n). What is the largest size n of the problem that can be solved in 1 second, assuming that the time required to solve the problem takes f(n) microseconds? How to start?

OpenStudy (anonymous):

So you want the value that will make lg(n) = 1 second or 1000000 microsecond

OpenStudy (anonymous):

I am still not sure how to proceed..

ganeshie8 (ganeshie8):

say n = 10^(10^6) then u wil get f(10^(10^6)) = log(10^(10^6)) us = 10^6 us = 1s

ganeshie8 (ganeshie8):

so the largest size of problem that can be solved in 1s is \(10^{1000000}\)

OpenStudy (anonymous):

10^6 us?? And why did you pick n = 10^(10^6)?

ganeshie8 (ganeshie8):

i just guessed, u can setup an equation aswell

ganeshie8 (ganeshie8):

we're given :- time required to solve the problem takes f(n) microseconds

ganeshie8 (ganeshie8):

so, if the problem size is n, then it takes f(n) microseconds

ganeshie8 (ganeshie8):

you need to solve the value of n, that makes f(n) = 1s

ganeshie8 (ganeshie8):

below eq'n :- 10^6 us = log(n)

ganeshie8 (ganeshie8):

remember that 10^6us = 1s

ganeshie8 (ganeshie8):

us = micro second

OpenStudy (anonymous):

Ah! So, if it is one minute instead of one second, then it will be 60 x 10^6 us = lg(n) ?

ganeshie8 (ganeshie8):

thats it, all bigOh problems are just primitive accounting problems lol

OpenStudy (anonymous):

Accounting is way easier than this. I just can't comprehend the problem. :(

ganeshie8 (ganeshie8):

did u get the phrase 'size of problem' ?

OpenStudy (anonymous):

No, not at all.

ganeshie8 (ganeshie8):

okay, suppose i give u \(10\) different numbers and ask you to sort them

ganeshie8 (ganeshie8):

we say, the size of this problem is \(10\)

ganeshie8 (ganeshie8):

and lets say, u take 1 minute to solve them.

ganeshie8 (ganeshie8):

next, suppose i give u 1000000 numbers and ask u to sort them this time u wil take more time, but how much time u wil take depends on wat algorithm u use to sovle them.

ganeshie8 (ganeshie8):

here problem size is 1000000

ganeshie8 (ganeshie8):

'problem size' and 'data size' both refer to same thing.

ganeshie8 (ganeshie8):

see if that seems plausible

OpenStudy (anonymous):

Ugh! Why is it not this: \[\frac{\lg n}{1}=\frac{1}{\lg n \times 10^{-6}}\]

ganeshie8 (ganeshie8):

dint get u ?

OpenStudy (anonymous):

Ah! Not that one.. No./Size of problem Time Given 1 lg(n) x 10^(-6) Find n 1 So, we get \[\frac{1}{n} = \frac{\lg n\times 10^{-6}}{1}\]

ganeshie8 (ganeshie8):

n = 1 is undefined i guess, cuz it wud make the time 0

ganeshie8 (ganeshie8):

consider the cases n > 1, and use bit common sense

ganeshie8 (ganeshie8):

if a thing is done in 1us, and we're given the relation that f(n) = log(n) give time for doing n things, finding how many things we can do in 1s should be a high-school grade problem. i dont see any point in you pondering over it :|

OpenStudy (anonymous):

" required to solve the problem takes f(n) microseconds" Isn't it lg(n) us instead of 1 us?

ganeshie8 (ganeshie8):

yes ur right

OpenStudy (anonymous):

Also, if 1 thing is done in lg(n) us, and n things done in 1s, then we can set the equation \[\frac{1}{\lg (n) \times 10^{-6}}=\frac{n}{1}\]Right? Wrong?

ganeshie8 (ganeshie8):

oops sorry, looks i was sounding bit offensive earlier

OpenStudy (anonymous):

It's ok...

ganeshie8 (ganeshie8):

10^6 us = log(n) whats the problem wid this equation ?

OpenStudy (anonymous):

I don't understand how you get this equation, and I don't understand why mine is not correct.

ganeshie8 (ganeshie8):

same is the case wid me, gimme a sec to digest ur equation

ganeshie8 (ganeshie8):

wid ur equaiton, wats the largest size you're getting ?

ganeshie8 (ganeshie8):

we may plug that number in reverse and see if it works or not

OpenStudy (anonymous):

Getting something weird with my equation. \[\frac{1}{\lg (n) \times 10^{-6}}=\frac{n}{1}\]\[n\lg (n) \times 10^{-6}=1\]\[\lg (n^n) =10^6\]\[n^n=10^{10^6}\] Don't know how to solve this thing..

ganeshie8 (ganeshie8):

humm ok, ima have dinner... wil get back and have a look at this again :)

ganeshie8 (ganeshie8):

read it like below :- Also, if \(\color{red}{n}\) things are done in lg(n) us, then find how many things can be done in 1s

OpenStudy (anonymous):

\[\frac{n}{\lg n \times 10^{-6}} = \frac{N}{1}\] N = Number of things done in one second.

ganeshie8 (ganeshie8):

still stuck on this ha ?

OpenStudy (anonymous):

Yes...

ganeshie8 (ganeshie8):

okay, i tried but im not getting ur equation, wat u trying to do wid that equation can u explain ?

ganeshie8 (ganeshie8):

we're given n things take log(n) micro seconds, and asked to find out how many things we can do in in 1 second so we simply sovle the equation log(n) = 10^6us

ganeshie8 (ganeshie8):

that gives the 'n' value corresponding to 10^6 us

OpenStudy (anonymous):

It's like, you need 4 dollars to buy a bar of chocolate, how many bars of chocolate you can buy when you have 120 dollars? We can have \[\frac{1}{4}=\frac{n}{120}\]

ganeshie8 (ganeshie8):

thats right, wat if it takes log(n) dollars to buy n chocolates ?

ganeshie8 (ganeshie8):

and you're asked to find how many u can buy wid a million dollars money ?

ganeshie8 (ganeshie8):

the price per chocolate is not fixed here, clearly it depends on how many u buy

OpenStudy (anonymous):

\[\frac{n}{\log n}=\frac{N}{1,000,000}\]?

ganeshie8 (ganeshie8):

its not a linear scaling, i dont think you can do it that way

ganeshie8 (ganeshie8):

first tell me this, are u convinced 10^(10^6) is the largest problem size yet ?

OpenStudy (anonymous):

How do you know if it is the largest?

ganeshie8 (ganeshie8):

plugin that number in the given function

OpenStudy (anonymous):

It works, but that does not guarantee there is no other (larger) solution, does it?

ganeshie8 (ganeshie8):

add 1 to that that, and plugin, wat do u get ?

ganeshie8 (ganeshie8):

u will get 1s + for any value larger than 10^(10^6) cuz log(n) is a strictly increasing function

ganeshie8 (ganeshie8):

but u want to complete job within 1s, so..

OpenStudy (anonymous):

Ah! Ok...

ganeshie8 (ganeshie8):

the question is begging us to solve f(n) = log(n) = 10^6 no more than that.

ganeshie8 (ganeshie8):

if it helps seeing this, let me plot the time Vs problem-size graph quick

ganeshie8 (ganeshie8):